OpenStudy (anonymous):

Find the general solution to these differential equations dy/dx=y-3

OpenStudy (anonymous):

@agent0smith

OpenStudy (anonymous):

the answer was \[y=Ce^x + 3\]

OpenStudy (agent0smith):

\[\frac{ dy}{ dx}=y-3 = 1(y-3)\] \[\frac{ dy }{ y-3 } = 1 dx\] Now integrate both sides, then you'll have to take the power of e on both sides.

OpenStudy (anonymous):

\[\frac{ 1 }{ y-3 }=x+c\]

OpenStudy (anonymous):

did i integrate it right?

OpenStudy (anonymous):

on the left hand side?

OpenStudy (agent0smith):

You might want to try a substitution on the right... let u = x-3, then the integral of 1/u is ln u

OpenStudy (anonymous):

wait nvm i got it lol

OpenStudy (anonymous):

can you help me with another one please? <3